1/2

William Blake,  William Blake,  William Blake,
The Tyger  Der Tiger  Le tigre
tradukita de Walter A. Aue  tradukita de Félix Rose

Tyger Tyger, burning bright, In the forests of the night; What immortal hand or eye, Could frame thy fearful symmetry?

Argiope-informo:

La aspekto de tiu ĉi versio estas barita.

Supozeble la aŭtoro decidis tion.

Tigre, tigre flamboyant d’ardeur

Dans les forêts de la nuit; Quel pouvoir immortel, inouï, Forma ta parfaite terreur?

In what distant deeps or skies Burnt the fire of thine eyes? On what wings dare he aspire? What the hand, dare sieze the fire?

Dans quels cieux, quelles profondeurs Brûlait le feu de ton regard? Quel vol osa donc vouloir? Et quelle main saisit ce feu?

And what shoulder and what art, Could twist the sinews of thy heart? And when thy heart began to beat, What dread hand? and what dread feet?

Et quelle puissance et quel art

Nouèrent les ligaments de ton coeur? Dès ses premiers battements, Quels pieds et griffes effrayant?

What the hammer? What the chain, In what furnace was thy brain? What the anvil? What dread grasp, Dare its deadly terrors clasp!

Quelle chaîne? Quel marteau?

Dans quelle fournaise était ton cerveau? Quelle enclume? Quel poing effarant Put gripper tes mortelles transes?

When the stars threw down their spears And water’d heaven with their tears: Did he smile his work to see? Did he who made the Lamb make thee?

Quand les astres jetant leurs lances Arrosèrent le ciel de pleurs, Sourit-Il en te contemplant, Lui qui fit l’agneau de douceur?

…  …  …

2/2

Tyger Tyger, burning bright, In the forests of the night; What immortal hand or eye, Dare frame thy fearful symmetry?

Tigre, tigre flamboyant d’ardeur Dans les forêts de la nuit; Quel pouvoir immortel, inouï, Osa former ta parfaite terreur?

Verkinto de tiu ĉi Angla poemo estas William Blake (∗1757-11-28 – †1827-08-12).

Arg-1048-2125 (2014-06-08 19:43:08)

Pri la poeto vidi la vikipediejon http: // eo. wikipedia. org/ wiki/ William_Blake .

Traduko de la Angla poemo “The Tyger ” de William Blake (∗1757-11-28 – †1827-08-12) en la Germanan de Walter A. Aue.

Arg-1048-2126 (2014-06-08 19:55:12)

Tiu ĉi poem-germanigo troviĝas en http: // myweb. dal. ca/ waue/ Trans/ Blake-Tiger. html .

Traduko de la Angla poemo “The Tyger ” de William Blake (∗1757-11-28 – †1827-08-12) en franca de Félix Rose.

Arg-1048-2129 (2014-06-09 11:37:35)

Tiu ĉi poem-traduko troviĝas en http: // donh. best. vwh. net/ Esperanto/ Literaturo/ Poezio/ tigro. html .

http://www.poezio.net/version?poem-id=1048&version-id=2125,2126,2129